# Energy Weighted Learning Progress Guided Interleaved Multi-Task Learning

### v1

@@ VI Conclusion
This paper introduced a biologically inspired interleaved multi-task learning framework that selects tasks dynamically based on learning progress and energy consumption. Motivated by how humans interleave tasks and regulate cognitive effort, our method, INTER-LP, prioritizes tasks that exhibit positive learning progress, while the energy-aware extension, INTER-LPE, further modulates selection by discounting the LP values, leading to a preference for the task with lower energy consumption. The model follows a shared encoder-decoder structure equipped with an attention mechanism that can learn multiple tasks with dynamic task switching. Our approach is evaluated on a set of action-effect prediction tasks in a simulated robotic environment.
@@
The experimental results demonstrate that LP-based task arbitration not only improves overall learning performance and convergence speed compared to baseline strategies like random interleaving, single-task learning, and blocked training, but also facilitates beneficial knowledge transfer across tasks. Moreover, the inclusion of energy-awareness in the task selection process enables a tunable trade-off between learning accuracy and resource efficiency, which is especially valuable in energy-constrained settings. Our ablation studies further validate the complementary roles of the attention mechanism and task flag bits in enhancing multi-task learning dynamics.

### v2

@@ 6 Conclusion
This paper introduced a biologically inspired interleaved multi-task learning framework that selects tasks dynamically based on learning progress and neural computational energy cost. Motivated by how humans interleave tasks and regulate cognitive effort, our method, IMTL-LP, prioritizes tasks that exhibit positive learning progress, while the energy-aware version, IMTL-EMLP, considers learning efficiency by discounting LP by the energy it takes to achieve it, shifting the preference to the tasks with lower energy consumption. The learning architecture developed follows a shared encoder-decoder structure equipped with bi-directional skill transfer mechanisms that can learn multiple tasks with dynamic task engagement.
@@
The developed framework is evaluated on a robotic manipulation scenario where the robot is tasked with learning the effects of its actions in different task environments. The experimental results demonstrate that LP-based task arbitration not only improves overall learning performance and convergence speed compared to baseline strategies like random interleaving, single-task learning, and blocked training, but also facilitates beneficial knowledge transfer across tasks. The results indicate that it is not only the training time allocated for the tasks that affect the learning performance, but also the task scheduling is critical, which is autonomously generated by our model. Moreover, the inclusion of energy-awareness in the task selection process enables a tunable trade-off between learning accuracy and resource efficiency, which is especially valuable in energy-constrained settings. Our ablation studies further validate the complementary roles of the attention mechanism and task flag bits in enhancing multi-task learning dynamics.
@@
Acknowledgements
This work was supported by the Japan Society for the Promotion of Science KAKENHI Grant Number JP23K24926. Additional support was provided by the project JPNP16007 commissioned by the New Energy and Industrial Technology Development Organization (NEDO) and Japan Society for the Promotion of Science KAKENHI Grant Number JP25H01236 and by the European Union under the INVERSE project (101136067).
Appendix A Task Settings and Hyperparameters
Module Layer In size Out size
$P^{\mathcal{T}_{i}}_{\text{state}}:\mathbb{R}^{d_{s}^{\mathcal{T}_{i}}}\rightarrow\mathbb{R}^{d_{s}}$ FC 9/9/18 4
$F:\mathbb{R}^{d_{s}}\rightarrow\mathbb{R}^{d_{h}}$ FC+ReLU 4 4
FC+ReLU 4 4
$f^{\mathcal{T}_{i}}:\mathbb{R}^{d_{h}}\rightarrow\mathbb{R}^{d_{r}}$ FC+ReLU 4 4
FC+ReLU 4 2
$MHA:\mathbb{R}^{d_{r}}\rightarrow\mathbb{R}^{d_{r}}$ MHA 2 2
$P^{\mathcal{T}_{i}}_{action}:\mathbb{R}^{d_{a}^{\mathcal{T}_{i}}}\rightarrow\mathbb{R}^{d_{a}}$ FC 8/8/12 1
$g^{\mathcal{T}_{i}}:\mathbb{R}^{d_{r}+d_{a}}\rightarrow\mathbb{R}^{d_{e}^{\mathcal{T}_{i}}}$ FC+ReLU 3 4
FC+ReLU 4 4
FC+ReLU 4 4
FC+ReLU 4 9/9/18
(a) Input and output dimensions for the single-task learning models.
Module Layer In size Out size
$P^{\mathcal{T}_{i}}_{\text{state}}:\mathbb{R}^{d_{s}^{\mathcal{T}_{i}}}\rightarrow\mathbb{R}^{d_{s}}$ FC 9/9/18 6
$F:\mathbb{R}^{d_{s}}\rightarrow\mathbb{R}^{d_{h}}$ FC+ReLU 6 6
FC+ReLU 6 4
$f^{\mathcal{T}_{i}}:\mathbb{R}^{d_{h}}\rightarrow\mathbb{R}^{d_{r}}$ FC+ReLU 4 4
FC+ReLU 4 2
$MHA:\mathbb{R}^{d_{r}+1}\rightarrow\mathbb{R}^{d_{r}+1}$ MHA 3 3
$P^{\mathcal{T}_{i}}_{action}:\mathbb{R}^{d_{a}^{\mathcal{T}_{i}}}\rightarrow\mathbb{R}^{d_{a}}$ FC 8/8/12 1
$g^{\mathcal{T}_{i}}:\mathbb{R}^{d_{r}+d_{a}+1}\rightarrow\mathbb{R}^{d_{e}^{\mathcal{T}_{i}}}$ FC+ReLU 4 4
FC+ReLU 4 4
FC+ReLU 4 4
FC+ReLU 4 9/9/18
(b) Input and output dimensions for the multi-task learning models.
Table 1: Network architecture details for both single and multi-task learning models. Number of neurons in $F$ in single-task learning is increased for the multi-task learning model, to ensure both architectures have the same number of parameters overall.
Appendix B IMTL-LP Task Selection Regimes
Figure 1: Different task selection regimes by our IMTL-LP model during the training. Counts are calculated with a sliding window $w=50$ and a step size $s=10$.